OpenStudy (anonymous):

Which calculation can be used to find the value of p in the equation p3 = 8?

OpenStudy (tkhunny):

That depends on what you mean. If it is \(p^{3}\) it's one thing. If you mean \(p\cdot 3\), it's another.

OpenStudy (anonymous):

Yes it's the first one p^3

OpenStudy (anonymous):

It will be p= cube root of 8 It's very easy

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \displaystyle x^\color{blue}{\rm n} =\color{#996633}{\rm a }}\) \(\LARGE\color{black}{ \displaystyle x=\sqrt[\color{blue}{\rm n} ]{\color{#996633}{\rm a }}}\) whether the \(\sqrt{\color{white}{\LARGE I}}\) will have a \(\pm\) in front of it, will depend on whether "n" is even or odd.

OpenStudy (anonymous):

@SolomonZelman there won't be +- sign in this case as we all know cube root of 8 is positive

OpenStudy (solomonzelman):

As if I said, it will definitely be positive .... (?)

OpenStudy (solomonzelman):

or neg...

OpenStudy (anonymous):

I'm trying to post the answers but i'm new so i'm having trouble.

OpenStudy (anonymous):

\[p=\sqrt[3]{8}\]

OpenStudy (anonymous):

\[p=3 \times \sqrt{8}\]

OpenStudy (anonymous):

p=2\[p=2\times \sqrt{8}\]

OpenStudy (anonymous):

\[p=\sqrt{8}\]

OpenStudy (anonymous):

@Samirah16 I didn't get ur reply. why did u write wrong answers?

OpenStudy (anonymous):

I was trying to write the answers